Reports 1-1 of 1 Clear search Modify search
Infrastructures (Air Conditioning)
soldani - 10:56 Wednesday 12 January 2022 (54459) Print this report
New set for DET Lab
As established yesterday by the meeting in this regard, this morning at around 8,30 (LT) the following new settings for the thermoregulation of the DET Lab were given:
Nominal mode
Frequency for supply fan: 40 Hz
Frequency for return fan: 17,5 Hz
Percentage of external air: 10%
In addition, the cylinder of the steam producer was replaced because it was exhausted
Search Help
×

Warning

×